OpenStudy (anonymous):

the area of a triangle is 21 square yards and its height is three yards. find the base???

OpenStudy (hoblos):

\[area = \frac{ base \times height }{ 2 }\] \[base = \frac{ 2 \times area }{ height }\]

OpenStudy (hoblos):

area= 21 and height=3 can you calculate the base ?

OpenStudy (anonymous):

yes thank you

OpenStudy (hoblos):

any time :)